I have a question that says the volumes of two spheres are 327pi mm cubed and 829pi mm cubed. What is the ratio of their surface areas?
Answer by mananth(16949)

volumes of two spheres are 327pi mm cubed and 829pi mm cubed. What is the ratio of their surface areas?
4/3 *pi^r^3= 327*pi
r^3= 3*327/4
r^3=981/4
r= cuberoot 981/4
6.26 mm
surface area = 4*pi*r^2
4*pi*(6.26)^2
.
4/3 *pi*r^3= 829 *pi
r^3= 2487/4
r=8.54 mm
surface area = 4*pi^(8.54)^2
Their ratio = 4*pi*(6.26)^2 / 4*pi^(8.54)^2
Ratio of surface areas = 0.73